can you help me solve 5y+7=4y-5 using the addition property??
Answer by lenny460(1073)

5y + 7 = 4y - 5
Subtract 4y from each side of the equation
5y + 7 - 4y = 4y - 5 - 4y
Combine like terms:
5y - 4y = 1y
4y - 4y = 0
Therefore:
1y + 7 = -5
Subtract 7 from each side of the equation
1y + 7 - 7 = -5 - 7
Combine like terms:
7 - 7 = 0
-5 - 7 = -12
1y = -12 or
y = -12


obuong3@aol.com